# Are Existing Road Design Guidelines Suitable for Autonomous Vehicles?

### v1

@@ 4.2 Evaluation
  • RQ4: Can $\mathtt{TrashFuzz}$ be used to test different ADSs?
  • RQ5: How difficult is it to carry out $\mathtt{TrashFuzz}$ attacks?
@@
RQ1: Can $\mathtt{TrashFuzz}$ produce valid scenarios? Given the specifications shown in Table I, we first investigate whether $\mathtt{TrashFuzz}$ can effectively generate valid scenarios that satisfy them. We compare $\mathtt{TrashFuzz}$ with randomly generated scenarios. Our evaluation involved analyzing 10,000 randomly generated scenarios, each with a fixed number of attack objects. Note that in these scenarios, we randomly place attack objects within a designated area, i.e., a 150m $\times$ 60m space in front of the AV. The results are shown in Table III. Here, the numbers (1 to 7) in the first row indicate the number of attack objects. We present the percentage of scenarios that satisfy the given specification. A higher percentage value indicates a greater number of valid scenarios. Rule2 and rule3 have been merged into one specification concerning the distance of fixed-position objects from the road. Likewise, rule4–6 are combined into one specification that evaluates the positioning of trees. This analysis demonstrates that a majority of the randomly generated scenarios are invalid with respect to public road design guidelines. Moreover, as the number of attack objects increases, the likelihood of scenarios deviating from these guidelines rises. This result indicates that most randomly generated scenarios are not ‘natural’. For instance, these objects are often placed directly on the road, a violation of $\mathtt{rule1}$ or $\mathtt{rule9}$ as outlined in Table I. As shown by the final row of Table III, $\mathtt{TrashFuzz}$ succeeds in generating a higher proportion of natural scenarios with respect to the placement guidelines in Table I.
TABLE IV: Violations of traffic laws while $\mathtt{TrashFuzz}$ operates under imposed restrictions
Traffic Laws LawBreaker $\mathtt{TrashFuzz}$ Law Concerns
@@ 0 Apollo - 7 9 5 7 7
The results are summarised in Table IV. Here, we present all the testable laws applicable to AVs within the platform. However, certain other relevant laws cannot be tested due to limitations in the available maps. For example, law 49 regulates that vehicles should not make U-turns at railway crossings, sharp bends, steep slopes, or tunnels. Unfortunately, the current maps of existing platforms (LGSVL and CARLA) do not include representations of these specific locations, rendering them untestable under the current conditions.
Within this table, the ‘LawBreaker’ and ‘$\mathtt{TrashFuzz}$’ columns indicate whether Apollo violated the specific traffic regulation listed in the first two columns. We mark $\surd$ for a traffic law $\phi$ if and only if $\pi\nvDash\phi$ happened for a scenario generated by LawBreaker or $\mathtt{TrashFuzz}$. We conducted multiple test runs, repeating the corresponding test cases at least three times to confirm and ensure the reproducibility of all issues. Note that the tools are violating laws using different mutation strategies: LawBreaker mutates the driving behaviour of other vehicles and pedestrians, whereas $\mathtt{TrashFuzz}$ instead mutates the placement of roadside objects subject to the roadside design guidelines (Table I). $\mathtt{TrashFuzz}$’s strategy is much more challenging due to the additional restrictions entailed.
As can be seen from the table, even while constrained to satisfy road design guidelines, $\mathtt{TrashFuzz}$ still manages to incite violations of a majority of the traffic laws. In comparison to LawBreaker, $\mathtt{TrashFuzz}$ covers all the traffic laws that LawBreaker violates. Additionally, $\mathtt{TrashFuzz}$ has the capability to activate an additional sub-law, specifically sub-law 7 of Article 51, which regulates that turning vehicles yield to straight-moving vehicles, whereas pedestrians and right-turning vehicles travelling in the opposite direction yield to left-turning vehicles. The activation of this supplementary sub-law is triggered by the presence of adversarial objects along the roadside. Note that $\mathtt{TrashFuzz}$ found an additional law violation compared to LawBreaker despite the restriction to modifying roadside object placements only.
We identified two broad categories of issues among the tests generated by $\mathtt{TrashFuzz}$, which we elaborate on below. These misjudgments stem from flaws in the perception system of the ADS. While the camera images and lidar points received by the ADS from the simulator appear normal, the system arrives at incorrect judgments. Note that we repeat the following issues at least three times to ensure their reproducibility.
Classification and Segmentation Errors. The first category of issues concerns misclassification of adversarial objects or malfunctions in the segmentation algorithm. For example, in some adversarial situations, trash bins may be erroneously classified as pedestrians or bicycles. In such cases, the ADS fails to identify these trash bins as stationary obstacles, which could potentially lead to collisions or other safety concerns (e.g. waiting for a ‘pedestrian’ to move when it is really a trash bin). Moreover, the deliberate placement of two trash bins in close proximity is intended to deceive the AV’s perception system into perceiving them as a single object such as a vehicle. This misclassification has the potential to confuse the vehicle’s decision-making algorithms, potentially causing hesitation or preventing the vehicle from taking appropriate action, particularly at intersections. Furthermore, when a trash bin is positioned lying down and adjacent to a bench, there are instances where the ADS fails to segment them separately, recognising them as a single unknown object. Note that these misclassifications are not limited to a single frame; rather, they last for a significant duration, propagating throughout the entire system and potentially inducing subsequent misbehavior, e.g. the ADS hesitating and failing to proceed as intended.
Overloaded Perception System. The second category of issues concerns the perception system of the ADS, which in general, is complex and consists of multiple algorithm and modules. In issues of this category, a few adversarial objects are strategically placed at specific positions along the ADS’s route to deny the service of the traffic light perception system of the ADS. In this situation, the traffic light perception system becomes overloaded, resulting in anomalous outputs. We observed two possible consequences of this situation in our tests. In the first, the roadside objects deceive the AV into perceiving the traffic light ahead as a red or yellow traffic light. This would cause the vehicle to interpret the intersection as a signal to stop indefinitely, regardless of the absence of an actual red traffic light or any other vehicles or pedestrians. Such a misperception can disrupt the normal flow of traffic and potentially lead to congestion or delays. In the second, we deceive the autonomous vehicle into perceiving the traffic light ahead as a green traffic light. Consequently, the detection system always indicates a green signal for the traffic light ahead, regardless of its actual state. This situation can lead to dangerous driving behaviours, such as rushing red lights, endangering the AV itself and others on the road.
@@
RQ3: Can our greedy gradient-based search algorithm cover more violations than the most comparable fuzzer? To answer this question, we do a baseline comparison of $\mathtt{TrashFuzz}$’s gradient-based search algorithm against the coverage-based fuzzing algorithm of LawBreaker [17], which is based on a genetic algorithm (GA) [43]. In particular, after decomposing traffic law specifications into violation goals (Section 3.2.2), we compare the coverage of the two tools in the sense of how many of those goals they are able to induce Apollo to satisfy.
@@
LawBreaker’s GA—the baseline in this experiment—begins by creating a generation of test scenarios and subsequently applies mutations and crossovers to produce the next generation. Intuitively, individuals with superior performance effectively ‘survive’ and contribute to the creation of the subsequent generations. In LawBreaker’s fuzzing algorithm, we initialise the population with 30 individuals and run the algorithm for 20 generations, resulting in a total of 620 test cases generated in a single execution. For $\mathtt{TrashFuzz}$, we configure it to allow a maximum of 620 queries, aligning with Lawbreaker’s fuzzing algorithm. Note that LawBreaker concentrates on altering the behavior of NPC vehicles and pedestrians on the road, while the focus of $\mathtt{TrashFuzz}$ is instead on modifying the placement of roadside objects. Therefore, to ensure a fair and meaningful comparison, we adopt identical test case encoding methodologies for both LawBreaker and $\mathtt{TrashFuzz}$, as detailed in Section 3.2.1. Additionally, we apply the same scenario restrictions outlined in Table I to LawBreaker’s fuzzing algorithm, to ensure our comparison solely assesses efficiency of the algorithms under equivalent conditions.
@@
The results are summarised in Table V. The overall traffic law specification $\phi$ we use is derived from LawBreaker [17], which contains 24 different Chinese traffic laws that can be decomposed into 82 possible violation goals ($\Theta(\phi)$). In the table, the ‘Num’ column denotes the quantity of adversarial objects involved in the respective attack scenario, ranging from 0 to 7. When the number is 0, it indicates that scenario is a common one without any attack. The ‘R1–4’ columns indicate how many of the 82 possible violation goals for $\phi$ are covered across four rounds of experimentation, where the average number is given in the last column. As shown, for both LawBreaker and $\mathtt{TrashFuzz}$, the coverage of violations increases proportionally with the number of adversarial objects. Furthermore, $\mathtt{TrashFuzz}$ surpasses LawBreaker in covering more violations within a limited number of queries.
@@
RQ4: Can $\mathtt{TrashFuzz}$ be used to test different ADSs? The attack carried out by $\mathtt{TrashFuzz}$ exploits the vulnerabilities in the perception algorithms of ADSs. These vulnerabilities are not limited to Apollo. We evaluated the perception of Autoware with the Carla simulator [44] as well. Here, we position adversarial objects at random locations along the AV’s path, varying their angles, and evaluate the perception results to determine whether the system ignores or misclassifies them.
In Table VI and Table VII, we present the results of the evaluations for Apollo (v7) and Autoware (Autoware.ai v1.14), respectively. The ‘Object’ column displays the name of each object, while the ‘Type’ column indicates whether the object is movable or of fixed-position. The ‘Vehicle?’, ‘Bicycle?’, ‘Pedestrian?’, and ‘Ignore?’ columns respectively check whether the perception of the ADS classifies the object as a vehicle, bicycle, pedestrian, or if it is ignored. It is worth noting that both Apollo and Autoware exhibit misclassifications for common objects along the road under certain circumstances. These misclassifications expose potential vulnerabilities in the perception system of these autonomous driving systems.

### v3

@@ IV-B Evaluation
  • RQ4: Do other ADS perception stacks exhibit similar object-induced misclassifications?
@@
RQ1: Can TrashFuzz produce valid scenarios? Given the specifications shown in Table I, we first investigate whether TrashFuzz can effectively generate valid scenarios that satisfy them. We compare TrashFuzz with randomly generated scenarios. Our evaluation involved analyzing 10,000 randomly generated scenarios, each with a fixed number of attack objects. Note that in these scenarios, we randomly place attack objects within a designated area, i.e., a 150m $\times$ 60m space in front of the AV. The results are shown in Table III. Here, the numbers (1 to 7) in the first row indicate the number of attack objects. We present the percentage of scenarios that satisfy the given specification. A higher percentage value indicates a greater number of valid scenarios. Rule2 and rule3 have been merged into one specification concerning the distance of fixed-position objects from the road. Likewise, rule4–6 are combined into one specification that evaluates the positioning of trees. This analysis demonstrates that a majority of the randomly generated scenarios are invalid with respect to public road design guidelines. Moreover, as the number of attack objects increases, the likelihood of scenarios deviating from these guidelines rises. This result indicates that most randomly generated scenarios are not natural. For instance, these objects are often placed directly on the road, a violation of $\mathtt{rule1}$ or $\mathtt{rule9}$ as outlined in Table I. As shown by the final row of Table III, TrashFuzz generates scenarios that satisfy the placement rules by construction (invalid mutations are discarded via lightweight constraint checks), yielding a much higher proportion of natural scenarios than random placement.
TABLE IV: Violations of traffic laws; note that TrashFuzz operates under natural scenario restrictions
Traffic Laws LawBreaker TrashFuzz Law Concerns
@@
The results are summarised in Table IV. Here, we present all the testable laws applicable to AVs within the platform. However, certain other relevant laws cannot be tested due to limitations in the available maps. For example, law 49 specifies that vehicles should not make U-turns at railway crossings, sharp bends, steep slopes, or tunnels. Unfortunately, the current maps of existing platforms (LGSVL and CARLA) do not include representations of these specific locations, rendering them untestable under the current conditions.
Within this table, the ‘LawBreaker’ and ‘TrashFuzz’ columns indicate whether Apollo violated the specific traffic regulation listed in the first two columns. We mark $\surd$ for a traffic law $\phi$ if and only if $\pi\nvDash\phi$ happened for a scenario generated by LawBreaker or TrashFuzz. We repeated each test case at least three times to ensure reproducibility and mitigate potential flakiness in simulation-based ADS testing [40]. Note that the tools are violating laws using different mutation strategies: LawBreaker mutates the driving behaviour of other vehicles and pedestrians, whereas TrashFuzz mutates only the placement of roadside objects subject to the roadside design guidelines (Table I). TrashFuzz’s strategy is thus much more challenging due to the additional restrictions.
As can be seen from the table, even while constrained to satisfy road design guidelines, TrashFuzz still manages to incite violations of a majority of the traffic laws. In comparison to LawBreaker, TrashFuzz covers all the traffic laws that LawBreaker violates. Additionally, TrashFuzz has the capability to activate an additional sub-law, specifically sub-law 7 of Article 51, which regulates that turning vehicles yield to straight-moving vehicles, whereas pedestrians and right-turning vehicles travelling in the opposite direction yield to left-turning vehicles. The activation of this supplementary sub-law is triggered by the presence of adversarial objects along the roadside. Note that TrashFuzz found an additional law violation compared to LawBreaker despite the restriction to modifying roadside object placements only.
We identified two broad categories of issues among the tests generated by TrashFuzz, which we elaborate on below. These misjudgments stem from flaws in the perception system of the ADS. While the camera images and lidar points received by the ADS from the simulator appear normal, the system arrives at incorrect judgments. Note that we repeat the following issues at least three times to ensure their reproducibility.
Classification and Segmentation Errors. The first category concerns misclassification of adversarial objects or failures in the segmentation algorithm. For instance, trash bins may be misclassified as pedestrians or bicycles, causing the ADS to treat them as moving agents rather than stationary obstacles (e.g., waiting for a ‘pedestrian’ that is actually a bin). Placing two bins in close proximity can also cause the perception system to interpret them as a single object such as a vehicle, potentially confusing downstream decision-making and causing hesitation or inappropriate behaviour at intersections. Similarly, when a bin lies adjacent to a bench, the ADS may fail to segment them separately and instead recognise a single unknown object. These errors persist across multiple frames, propagating through the system and inducing misbehaviour such as hesitation or failure to proceed.
Overloaded Perception System. The second category of issues concerns the perception system of the ADS, which in general, is complex and consists of multiple algorithms and modules. In issues of this category, a few adversarial objects are strategically placed at specific positions along the ADS’s route to deny the service of the traffic light perception system of the ADS. In this situation, the traffic light perception system becomes overloaded, resulting in anomalous outputs. We observed two possible consequences of this situation in our tests. In the first, the roadside objects deceive the AV into perceiving the traffic light ahead as a red or yellow traffic light. This would cause the vehicle to interpret the intersection as a signal to stop indefinitely, regardless of the absence of an actual red traffic light or any other vehicles or pedestrians. Such a misperception can disrupt the normal flow of traffic and potentially lead to congestion or delays. In the second, we deceive the autonomous vehicle into perceiving the traffic light ahead as a green traffic light. Consequently, the detection system always indicates a green signal for the traffic light ahead, regardless of its actual state. This situation can lead to dangerous driving behaviours, such as rushing red lights, endangering the AV itself and others on the road.
@@ 0 Apollo - 7 9 5 7 7
RQ3: Can our greedy gradient-based search algorithm cover more violations than the most comparable fuzzer? To answer this question, we do a baseline comparison of TrashFuzz’s gradient-based search algorithm against the coverage-based fuzzing algorithm of LawBreaker [50], which is based on a genetic algorithm (GA) [34]. In particular, after decomposing traffic law specifications into violation goals (Section III-B2), we compare the coverage of the two tools in the sense of how many of those goals they are able to induce Apollo to satisfy.
@@
LawBreaker’s GA—the baseline in this experiment—begins by creating a generation of test scenarios and subsequently applies mutations and crossovers to produce the next generation. Intuitively, individuals with superior performance effectively survive and contribute to the creation of the subsequent generations. In LawBreaker’s fuzzing algorithm, we initialise the population with 30 individuals and run the algorithm for 20 generations, resulting in a total of 620 test cases generated in a single execution. For TrashFuzz, we configure it to allow a maximum of 620 queries, aligning with LawBreaker’s fuzzing algorithm. Note that LawBreaker concentrates on altering the behavior of NPC vehicles and pedestrians on the road, while TrashFuzz focuses on modifying the placement of roadside objects. Therefore, to ensure a fair and meaningful comparison, we adopt identical test case encoding methodologies for both LawBreaker and TrashFuzz, as detailed in Section III-B1. Additionally, we apply the same scenario restrictions outlined in Table I to LawBreaker’s fuzzing algorithm, to ensure our comparison solely assesses efficiency of the algorithms under equivalent conditions.
@@
The results are summarised in Table V. The overall traffic law specification $\phi$ we use is derived from LawBreaker [50], which contains 24 different Chinese traffic laws that can be decomposed into 82 possible violation goals ($\Theta(\phi)$). In the table, the ‘Num’ column denotes the quantity of adversarial objects involved in the respective attack scenario, ranging from 0 to 7. When the number is 0, it indicates that scenario is a common one without any attack. The ‘R1–4’ columns indicate how many of the 82 possible violation goals for $\phi$ are covered across four rounds of experimentation, where the average number is given in the last column. As shown, for both LawBreaker and TrashFuzz, the coverage of violations increases proportionally with the number of adversarial objects. Furthermore, TrashFuzz surpasses LawBreaker in covering more violations within a limited number of queries.
@@
RQ4: Do other ADS perception stacks exhibit similar object-induced misclassifications? The attack carried out by TrashFuzz exploits vulnerabilities in ADS perception algorithms, which are not limited to Apollo. Fully integrating TrashFuzz with other ADS stacks is non-trivial due to architectural differences (e.g., Autoware’s ROS-based microservice architecture). Therefore, rather than running the full TrashFuzz optimisation loop, we conduct a preliminary perception-level probe using Autoware with the CARLA simulator [14]. We place roadside objects at random locations along the AV’s path, vary their orientations, and evaluate whether the perception system ignores or misclassifies them.
In Table VII and Table VII, we present the results of the evaluations for Apollo (v7) and Autoware (Autoware.ai v1.14), respectively. The ‘Object’ column displays the name of each object, while the ‘Type’ column indicates whether the object is movable or of fixed-position. The ‘Vehicle?’, ‘Bicycle?’, ‘Pedestrian?’, and ‘Ignore?’ columns respectively check whether the perception of the ADS classifies the object as a vehicle, bicycle, pedestrian, or if it is ignored. It is worth noting that both Apollo and Autoware exhibit misclassifications for common objects along the road under certain circumstances. These misclassifications expose potential vulnerabilities in the perception system of these autonomous driving systems.
Although our evaluation focuses on Apollo, we conducted these preliminary experiments with Autoware to explore TrashFuzz’s generalisability. These showed that adversarial layouts generated by TrashFuzz can also induce classification errors in Autoware, suggesting that the underlying attack strategies may transfer across different ADSs. However, full pipeline integration is currently limited by Autoware’s ROS-based microservice architecture, which complicates the incorporation of our constraint-based scenario generation and STL-guided optimisation loop. Future work will therefore focus on adapting TrashFuzz for broader ADS compatibility by: (i) developing ROS-compatible modules for scenario evaluation and STL violation detection, (ii) introducing an abstraction layer between the TrashFuzz optimiser and modular perception outputs, and (iii) deploying the framework on additional platforms such as CARLA with Autoware.Universe. These steps would extend the applicability of TrashFuzz and enable robust cross-platform adversarial testing.
V Discussion
V-A Threats to Validity
Simulation-based testing introduces validity concerns. ADSs require significant computational resources, and insufficient resources may add latency. To mitigate this, we ran Apollo and the simulator on separate machines connected via Ethernet, ensured memory cleanliness, and avoided background tasks. Under these conditions, Apollo’s modules maintained an average latency below 0.1s. We also reproduced each identified issue at least three times. While environment-induced effects cannot be fully ruled out, the surfaced behaviours remain valuable for understanding and improving ADS robustness.
V-B Practicality of TrashFuzz Attacks
We briefly discuss the practical assumptions required to realise attacks such as TrashFuzz in comparison with existing approaches. Table VIII summarises the prerequisites of several representative attacks in the literature.
Within the table, the categories ‘Patch?’, ‘Shape?’, ‘Placement?’, and ‘Special?’ indicate whether the attack requires a specialised adversarial patch, modifies the shape of an object, relies on carefully chosen placement of physical objects, or requires specialised equipment unlikely to be found on the roadside (e.g., commercial drones, digital LiDAR spoofers, or projectors). The columns ‘White-B?’, ‘Grey-B?’, and ‘Black-B?’ indicate the assumed level of access to the ADS. A checkmark ($\surd$) denotes that a particular prerequisite is required. For example, the MSF-Attack [5] modifies the shape of roadside objects (e.g., traffic cones) to make them invisible to camera and LiDAR detectors, and therefore requires both object shape manipulation and controlled placement.